Diamond Cleats Co. manufactures cleats for baseball shoes. It has outlined the following overhead cost drivers:

Overhead Costs Pool Cost Driver Overhead Cost
Budgeted Level
for Cost Driver
Budgeted
Quality control # of inspections $ 87,100 1,300
Machine time Machine hours 255,120 1,063
Materials handling # of Batches 1,650 55
Miscellaneous overhead cost Direct labor hours 64,000 6,700

Diamond Cleats Co. has an order for cleats that has the following production requirements:

Number of inspections 390
Number of machine hours 230
Number of batches 8
Direct labor hours 920

Using activity-based costing, applied quality control factory overhead for the baseball cleat order is:

Multiple Choice

  • $28,550.

  • $30,320.

  • $26,130.

  • $21,250.

  • $19,700.
Answer #1

Answer : $ 26,130

Quality Control Rate = Estimated cost Driver / Estimated Cost Driver

Quality Control Rate = 87,100 /1300 = $67

Applied Quality Control = Inspection * Rate = 390*67 = $ 26,130 (Answer)